if A:c=1:3 and b:c=5:7,find a:b and a:b:c.
Answer by RAY100(1637)

a/b=a/c*c/b = 1/3 * 7/5 = 7/15
.
a/b/c,, a/b=7/15,,,and b/c =5/7 or 15/21,,therefore a/b/c = 7/15/21
.
look at it as a/b and b/c ,,,7:15,,,and 15:21,,,or 7:15:21